OpenStudy (anonymous):

for the function f(x)=(8-x^3)/(x^2-4) evaluate the function the values of x compute 2.5

mathslover (mathslover):

\[\large{f(x) = \frac{(8-x^3)}{(x^2-4)}}\] \[\large{f(x) = \frac{(2^3-x^3)}{(x^2-2^2)}}\] \[\large{f(x) =\frac{(2+x)(4+x^2-2x)}{(x-2)(x+2)}}\]

OpenStudy (campbell_st):

well factorise the numerator and demonimator 1st the numerator is the difference of 2 cubes... the denominator is the difference of 2 square

mathslover (mathslover):

Just like I did above :)